Measurements of height, weight, body circumferences, and body fat are called

Sonography.
Electrocardiography.
Echocardiography.
Anthropometry.

Which of the following are substances in plant foods that are not digested in the stomach or small intestine?

Dietary fiber
Dextrose
Simple sugars
Disaccharides

An infant receiving a well-child exam has his length, weight and head circumference measured.  This part of the examination is the:

Anthropometric assessment
Physical examination
Biochemical evaluation
Nutrition assessment

Which term describes failing health that results from a long-standing dietary intake that does not meet nutritional needs?

Inferior nutrition
Subclinical nutrition
Balanced nutrition
Undernutrition

Which of the following measures does not assess nutritional status?

Clinical evaluation
Dietary analysis
Biochemical evaluation
Hair analysis

Subclinical nutrition status indicates a person:

Will not exhibit symptoms of nutrient deficiency
Is very malnourished and suffering malnutrition symptoms
Is well nourished
Is at risk for heart disease from excess calorie intake

One serving of Raisin Bran cereal contains 50 percent of the 18 mg Daily Value for iron.  How much iron will one serving of your cereal provide?

36 mg
15 mg
26 mg
9 mg

All of the following are key recommendations from the US Dietary Guidelines for Americans except:

If consumed, limit alcohol to fewer than one drink per day for women and fewer than two for men
Choose nutrient-dense foods
Eliminate added sugars
Half of grain intake should be from whole grains
Answer #1

ANSWER

* Measurement of height,weight,body circumference and body fat

is known as Anthropometry.

* Dietry fibers present in plant foods are not digested in stomach or small

intestine.

* Examination of length,weight,head circumference in well child exam is a

   part of Anthropometric assessment.

* Undernutrition is the term which describes fall health that result of long standing

dietry intake that does not meet nutritional needs.

* Measures which does not assess nutritional status is Hair analysis.

* Subclinical nutritional status indicates a person will not exhibit symptoms of nutritional

   deficiency.

* Calculation of amount iron in one serving of cereal.

            One serving = 50% 0f 18 mg daily value of iron.

                        50% = 50 / 100

                                = 50/100 * 18

                                  = 0.5 * 18

                                = 9 mg.

* All of the above are key recommendation from the US dietary guidelines for

Americans Except Choose nutrient-dense foods.